OpenStudy (anonymous):

find the solution of :

OpenStudy (anonymous):

\[\sqrt{x^{2} - 3x - 10} - 2\sqrt{x^{2} - 3x - 9} + \sqrt{x^{2} - 3x - 8} < 0\]

OpenStudy (anonymous):

i take u = x^2 - 3x - 10, give sqrt(u) - 2 sqrt(u + 1) + sqrt(u + 2) < 0 sqrt(u) + sqrt(u + 2) < 2sqrt(u+1) 2u + 2 + 2sqrt(u^2 + 2u) < 4u + 4 2 sqrt(u^2 + 2u) < 2u + 2 4(u^2 + 2u) < 4u^2 + 8u + 16 o < 16 i was missing of u :D

OpenStudy (rational):

Looks good !

OpenStudy (anonymous):

i saw an answer through wolfram, but how ? there is no explanation

OpenStudy (anonymous):

the solution is x<= -2 or x>= 5 (wolfram's version) but i need a way

myininaya (myininaya):

the 4th line can only happen if u>0 and u+2>0 the intersection of those sets are u>0 So we want u>0 which means x^2-3x-10>0 <--is what you need to solve

myininaya (myininaya):

solving this will make your last line true you know that 0<16

OpenStudy (anonymous):

actually, i know just square both sides any times to get u and the rule that the domain in sqrt must be >= 0 sqrt(u) + sqrt(u + 2) < 2sqrt(u+1) how about the right side ?

OpenStudy (rational):

x^2-3x-10>=0

myininaya (myininaya):

I don't understand the question you are asking @tanjung

myininaya (myininaya):

and yes \[\sqrt{A} \cdot \sqrt{B}=\sqrt{A \cdot B } \text{ \iff } A \ge 0 \text{ and } B \ge 0\]

myininaya (myininaya):

you did this in your fourth line

OpenStudy (anonymous):

2 sqrt(u^2 + 2u) < 2u + 2 is that ?

OpenStudy (rational):

\(\sqrt{u} - 2 \sqrt{u + 1} + \sqrt{u + 2} < 0\) yields \(0 < 16\) there is no mistake in your work but you also need to look at the domain of \(u\) : \[u\ge 0 ,~~~u+1\ge 0,~~~u+2\ge 0\]

myininaya (myininaya):

right

myininaya (myininaya):

intersection of all those sets mentioned by rational is u>=0

OpenStudy (anonymous):

so, we dont need find u then solve for x ?

myininaya (myininaya):

your inequality 0<16 is true for u in the domain of the original equation

myininaya (myininaya):

so 0<16 happens when u>=0 replace u with what you set it equal to and solve that inequality for x

OpenStudy (rational):

we do need to find \(u\) then solve for \(x\). find \(u\) from below set of inequalities : \[0\lt 16,~~u\ge 0 ,~~~u+1\ge 0,~~~u+2\ge 0\] Notice the first inequality is an useless inequality

myininaya (myininaya):

I should have included u+1>=0 earlier and also look at the original equation I for some reason just looked at the sqrt(AB)=sqrt(A)*sqrt(B) part :p

myininaya (myininaya):

but anyways the intersection of all those sets in terms of u is still the same

OpenStudy (anonymous):

oh, ok i think i can understand now. thanks for you both @myininaya and @rational :D